OpenStudy (dj3strella):

Can someone help me please I for got how to solve this problem. Solve for m. n 1/2 (m+p)

OpenStudy (anonymous):

Show me your choices

OpenStudy (dj3strella):

the choices are a.) m =1/2p-n b.) m =2n-p c.) m =2p+n d.) m=1/2n-p @Shay17

OpenStudy (anonymous):

yea so D.